MIT/CXC Group Meeting, 27 July 1999

Chandra Status (Claude; Peter): ACIS vented, cooled. Extra two days needed to acheive orbit, perigee 750 km lower than planned, but OK. Everything is going very well.

Duty Scientist shift report (Dave D.): VCDU (spacecraft clock) blocks duplicated, confuses processing. Being fixed by GOT (Ground Operations Team). HRC pulse data look good; aspect FID-light data processed. Level 0 processing working, Level 0.5 and 1 working.

Grating ARF/RMF work (Dave H.): making nominal prototype files available. DPH --- grating ARFs (effective area), for on-axis, no dither --- need to convert to XSPEC-style formats. RMF - Chris B. producing single-order RMFs.

Extended sources: --- big gap is in vizualization/selection of regions, propagation of region info to tools and files. Need good simulation, Level 1.5 product for experimentation, and "blind" analysis, in order to determine what new products or techniques are needed.

HETG insertion issues (Dan D.): want to watch Capella for 5ks observation while inserted. Occurs around 21 August.

ISIS "Rehearsal" (Dan D.): use Capella simulation, process to L2 (binned spectrum), give to someone for analysis.

MARX Release (Mike W.): 2.22 release next week. What's next? Ideas: CCD model (for pileup); ACIS modes; HRC sub-atomic model.

Misc. infrastructure (Dave H.): poster printer - should we get one? room for students --- renovations in the queue; 72GB of disks - --> --ordered; CPUs: need more for student/guests.

ACTIONS: (explicit or implied)

David P. Huenemoerder